3x^2-5x+3y^2-8=0
Write the equation in standard form. Identify which type of conic section it is: circle, parabola, hyperbola, or ellipse. Graph and label the appropraite parts. Show work.

Answer by lwsshak3(11628)

3x^2-5x+3y^2-8=0
Write the equation in standard form. Identify which type of conic section it is: circle, parabola, hyperbola, or ellipse. Graph and label the appropraite parts. Show work.
..
3x^2-5x+3y^2-8=0
3(x^2-5x/3+25/36)+3y^2=8+25/12=96/12+25/12=121/12
3(x-5/6)^2+3(y+0)^2=121/12
divide both sides by 3
(x-5/6)^2+(y+0)^2=121/36
(x-5/6)^2+y^2=121/36
This is a circle with center at (5/6,0) and radius=sqrt(121/36)=1.833
see graph below as a visual check on the answer:
..
y=(121/36-(x-5/6)^2)^.5

..